# LogLLM: Log-based Anomaly Detection Using Large Language Models

### v3

@@ II-B LLMs for Log-based Anomaly Detection
Figure 2: Session window.
Figure 3: Fixed window.
@@ III Preliminaries
The log messages can be grouped into log sequences (i.e., series of log messages that record specific execution flows) based on sessions or fixed/sliding windows [48]. Session window partitioning groups log messages according to their session IDs, thereby generating sequences that include the log messages within each session. For example, Fig. 2 illustrates the HDFS [49] logs undergoing the session window grouping process, where the block_id serves as the session ID. In contrast, fixed/sliding window partitioning groups log messages based on a fixed size (i.e., window size), which can be defined by either the time span or the number of log messages. This method creates sequences that capture snapshots of system log messages over time. For example, Fig. 3 illustrates the BGL [50] logs undergoing the fixed window grouping process, with a window size of 2 messages and a step size of 2 messages.
@@ IV-A Preprocessing
Thanks to the structured log generation process, the textual format of parameters representing specific objects can be easily identified using regular expressions [53]. Consequently, we replace each variable parameter, such as account, directory path, and IP address, with ’<*>’. Despite its simplicity, this technique offers significant performance advantages. Compared with log parsers, this preprocessing technique is more effective and does not require training.
@@ IV-B3 Llama
The first component introduces the log sequence, such as "Below is a sequence of system log messages:". The second component comprises the token embeddings $E$ output by the projector. The third component queries whether the sequence is anomalous, asking, for instance, ". Is this sequence normal or anomalous?". The first and third components are fed into the Llama tokenizer and Llama embedding layer sequentially, producing $E_{1}\in\mathbb{R}^{A\times d_{Llama}}$ and $E_{3}\in\mathbb{R}^{Q\times d_{Llama}}$, where $A$ and $Q$ are the number of tokens produced by tokenizing the first and third components, respectively. Then, the token embeddings of the three components are concatenated, represented as $[E_{1}||E||E_{3}]\in\mathbb{R}^{(A+N+Q)\times d_{Llama}}$ and fed into Llama. We utilize Llama 3 8B  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B for this process, where $d_{Llama}$ is 4096.
@@ IV-C1 Minority Class Oversampling
LogLLM is a supervised anomaly detection method, which means it needs labeled normal and anomalous samples for training. However, supervised anomaly detection methods often face the challenge of data imbalance, which can lead to biased model training. In an anomaly detection task, there are only two classes: normal and anomalous, and the number of instances in each class is uncertain. To cope with data imbalance, we oversample the class with fewer samples, ensuring that the proportion of the minority class is no less than $\beta$. Formally, let the the proportion of the minority class is $\alpha$ and $\alpha<\beta$, and the total number of samples is $Sample\_num$. To achieve a proportion of $\beta$ for the minority class, it will be oversampled to the following quantity:
@@ IV-C2 Training Objective
Our objective is to train the deep model to predict whether a given log sequence is normal or anomalous. We fine-tune the model to respond appropriately: if the sequence is anomalous, it outputs ’The sequence is anomalous’; if normal, it outputs ’The sequence is normal’. We utilize cross-entropy loss [54] as our loss function.
@@ IV-C3 Training Procedure
Stage 1. Fine-tuning Llama to capture the answer template: The first stage involves fine-tuning Llama to capture the answer template. Specifically, we train Llama to respond to the prompt ’Is this sequence normal or anomalous?’ with ’The sequence is anomalous/normal’. This stage requires only a few data samples.
@@ V Experiments
  • RQ3: Is each stage in the three-stage training procedure necessary?
  • RQ4: Is oversampling of the minority class necessary? Additionally, how does the hyperparameter $\beta$ influence the performance of LogLLM?
LogLLM is coded in Python, and the source code is available at https://anonymous.4open.science/r/LogLLM-B7EA.
@@ V-D Dataset
HDFS (Hadoop Distributed File System) dataset [49] is generated by running Hadoop-based mapreduce jobs on over 200 Amazon EC2 nodes and contains a total of 11,175,629 log messages. These log messages are grouped into different log windows based on their block_id, which reflect program executions in the HDFS. Among these, 16,838 blocks (2.93%) indicate system anomalies.
@@
In the context of HDFS, we adopt a session window strategy, which involves grouping log messages into sequences based on the block_id present in each log message. Each session is labeled using ground truth. For other datasets, including BGL, Thunderbird, and Liberty, we utilize a fixed window strategy to group log messages, with a window size of 100 messages and a step size of 100 messages. A log sequence is deemed anomalous if it contains at least one anomalous log message according to the ground truth.
@@ V-E Performance Evaluation (RQ1)
Effect of labeled anomalies: As illustrated in Table II, in contrast to methods such as DeepLog, LogAnomaly, FastLogAD, LogBERT, and RAPID, which require clean datasets devoid of anomalies to build anomaly detection models, methods like PLELog, LogRobust, CNN, NeuralLog, and LogLLM demonstrate superior performance. These models are trained using not only normal samples but also labeled anomalies. For instance, these five methods achieve an average F1-score above 0.771 across four datasets, whereas others that do not utilize labeled anomalies perform poorly, with an average F1-score below 0.602 across four datasets. This demonstrates that incorporating labeled anomalies can provide a significant advantage to anomaly detection methods.
@@
V-F Effects of Different Preprocessing Techniques (RQ2)
@@
We evaluate the effectiveness of the different preprocessing techniques. The results are shown in Table IV. In this table, ’Raw’ indicates that the content of log messages is not preprocessed and is directly input into the proposed deep model. ’Template’ indicates that sequences of log templates produced by Drain [51], a log parser, are used as input for the proposed deep model. ’Template ID’ signifies that the IDs of log templates, obtained by Drain, are simply encoded into numeric vectors using an embedding layer instead of BERT. The preprocessing technique ’Template ID’ renders the model unable to capture the semantic information within log messages. Notably, the parser Drain is applied to the entire dataset, rather than only the training dataset, to avoid performance degradation due to the OOV problem. ’RE’ indicates that regular expressions, as introduced in Section IV-A, are used for preprocessing log messages.
As anticipated, the preprocessing technique ’RE’ yields the highest F1-score across all datasets. Conversely, the preprocessing technique ’Template ID’ consistently results in the lowest F1-score across all datasets, averaging 36.8% lower than that of ’RE’. This can be attributed to the fact that ’Template ID’ hinders the model’s ability to capture the semantic information within log messages, thereby impairing its capability to detect anomalies from a natural language perspective. The preprocessing techniques ’Raw’ and ’Template’ result in relatively good performance, but their F1-scores are still 6.4% and 5.7% lower than that of ’RE’, respectively. For the preprocessing technique ’Raw’, the variable parts (parameters that carry dynamic runtime information) within the content of each log message have little influence on anomaly detection. However, due to their high randomness, they can confuse the model, making it difficult to discern anomalies. For the preprocessing technique ’Template’, the parser is not always reliable, sometimes incorrectly removing the constant parts or retaining the variable parts, which can lead to information loss or confusion for the model, making it difficult to discern anomalies.
V-G Ablation Study of the Training Procedure (RQ3)
@@
We investigate the effect of each training procedure through an ablation study. The results are presented in Table V, where ’W/O’ denotes ’without’. We have the following observations:
Skipping any training stage results in a decrease in the F1-score across all datasets, demonstrating the effectiveness of our three-stage training procedure. Training without stage 1 leads to the worst performance, with the F1-score averaged across all datasets decreasing by as much as 29.7%. This demonstrates that fine-tuning Llama to capture the answer template (Stage 1) is essential before training the embedder of log messages (Stage 2). Without this stage, the embedder may be misdirected, resulting in incorrect semantic capture of log messages and model failure. Training without stage 3 yields relatively poor performance, with an average F1-score decrease of 10.5%. This indicates that sequentially fine-tuning Llama and training the embedder alone is insufficient for the model to capture anomalous patterns; cohesive fine-tuning of the entire model is essential. Training without stages 2 and 1&2 (only adopting training stage 3: fine-tuning the entire model) results in acceptable performance, with average F1-score decreases of 2.5% and 3.2%. This demonstrates that individually training the embedder (BERT and projector) before fine-tuning the entire model can also enhance performance. This stage allows the embedder to generate better semantic vectors of log messages for Llama to discern anomalies.
@@
V-H Impact of Minority Class Oversampling (RQ4)
Note that normal and anomalous samples in the training dataset are imbalanced, as shown in Table I. For the HDFS, BGL, and Thunderbird datasets, normal samples outnumber anomalous samples. Conversely, in the Liberty dataset, anomalous samples exceed normal samples. The hyper-parameter $\beta$ controls the proportion of the minority class by oversampling to address data imbalance problem, as described in Section IV-C1. In this section, we investigate the impact of $\beta$ by varying its value. Fig. 5 illustrates the performance of LogLLM on the four datasets under different magnitudes of $\beta$. When $\beta=0$, the samples are not oversampled; instead, the original datasets are utilized directly for training.
@@ V-H Impact of Minority Class Oversampling (RQ4)
As illustrated in Fig. 5c, the changing pattern of the F1-score is basically the same across all datasets. The F1-score increases and then decreases as $\beta$ increases. However, the LogLLM seems not to be sensitive to $\beta$; when $\beta$ is between 10% and 80%, the variation in the F1-score is no more than 0.07. Thanks to the substantial semantic knowledge embedded in LLMs, a trained model can effectively learn anomalous patterns and detect anomalies, even when the minority class constitutes only 10% of the dataset. In comparison to the BGL and Thunderbird datasets, the F1-score for the HDFS and Liberty datasets shows minimal variation with respect to $\beta$. This consistency can be attributed to the clear anomalous patterns present, allowing LogLLM to be easily trained to detect anomalies across various $\beta$ values. However, LogLLM appears unable to effectively handle extremely imbalanced scenarios. For instance, in the Thunderbird dataset, anomalies constitute only 1.05% of the samples, causing the trained model to be biased and classify all samples as normal. As a result, precision, recall, and F1-score are all equal to 0. Consequently, minority class oversampling is sometimes essential.

### v4

@@ III Preliminaries
(a) Session window
(b) Sliding window
Figure 2: Illustrative examples of log message partitioning.
The log messages can be grouped into log sequences (i.e., series of log messages that record specific execution flows) based on session or fixed/sliding windows [48]. Session window partitioning groups log messages according to their session IDs, thereby generating sequences that include the log messages within each session. For example, Fig. 2a illustrates the HDFS [49] logs undergoing the session window grouping process, where the block_id serves as the session ID. In contrast, fixed/sliding window partitioning groups log messages based on a fixed size (window size), which can be defined by either the time span or the number of log messages. This method creates sequences that capture snapshots of system log messages over time. For example, Fig. 2b illustrates the BGL [50] logs undergoing the sliding window grouping process, with a window size of 2 messages and a step size of 2 messages.
@@ IV-A Preprocessing
Thanks to the structured log generation process, the textual format of parameters representing specific objects can be easily identified using regular expressions [53]. Consequently, we replace each variable parameter, such as account, directory path, and IP address, with ‘<*>’. Despite its simplicity, this technique offers significant performance advantages. Compared with log parsers, this preprocessing technique is more effective and does not require training.
@@ IV-B3 Llama
The first component introduces the log sequence, such as "Below is a sequence of system log messages:". The second component comprises the token embeddings $E$ output by the projector. The third component queries whether the sequence is anomalous, asking, for instance, ". Is this sequence normal or anomalous?". The first and third components are fed into the Llama tokenizer and Llama embedding layer sequentially, producing $E_{1}\in\mathbb{R}^{A\times d_{Llama}}$ and $E_{3}\in\mathbb{R}^{Q\times d_{Llama}}$, where $A$ and $Q$ are the number of tokens produced by tokenizing the first and third components, respectively. Then, the token embeddings of the three components are concatenated, represented as $[E_{1}||E||E_{3}]\in\mathbb{R}^{(A+N+Q)\times d_{Llama}}$ and fed into Llama. We utilize Llama 3 8B model [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B, which consists of 32 transformer decoder-like layers and contains 8 billion parameters. Here, $d_{Llama}$ is 4096.
@@ IV-C1 Minority Class Oversampling
LogLLM is a supervised anomaly detection method, which means it needs labeled normal and anomalous samples for training. However, supervised anomaly detection methods often face the challenge of data imbalance, which can lead to biased model training. In an anomaly detection task, there are only two classes: normal and anomalous, and the number of instances in each class is uncertain. To cope with data imbalance, we oversample the class with fewer samples, ensuring that the proportion of the minority class is no less than $\beta$. Formally, let the proportion of the minority class be $\alpha$ and $\alpha<\beta$, and the total number of samples be $Sample\_num$. To achieve a proportion of $\beta$ for the minority class, it will be oversampled to the following quantity:
@@ IV-C2 Training Objective
Our objective is to train the deep model to predict whether a given log sequence is normal or anomalous. We fine-tune the model to respond appropriately: if the sequence is anomalous, it outputs ‘The sequence is anomalous.’; if normal, it outputs ‘The sequence is normal.’. We utilize cross-entropy loss [54] as our loss function.
@@ IV-C3 Training Procedure
Stage 1. Fine-tuning Llama to capture the answer template: The first stage involves fine-tuning Llama to capture the answer template. Specifically, we train Llama to respond to the prompt ‘Is this sequence normal or anomalous?’ with ‘The sequence is anomalous/normal.’. This stage requires only a few data samples.
@@ V Experiments
  • RQ3: How does each stage of the three-stage training process influence the performance of LogLLM?
  • RQ4: How do different levels of minority class oversampling, determined by the hyperparameter $\beta$, affect the performance of LogLLM?
LogLLM is coded in Python, and the source code is available at https://github.com/guanwei49/LogLLM.
@@ V-D Dataset
HDFS (Hadoop Distributed File System) dataset [49] is generated by running Hadoop-based mapreduce jobs on over 200 Amazon EC2 nodes and contains a total of 11,175,629 log messages. These log messages are grouped into different log windows based on their block_id, which reflect program executions in the HDFS, resulting in 575,061 blocks. Among these, 16,838 blocks (2.93%) indicate system anomalies.
@@
In the context of HDFS, we adopt a session window strategy, which involves grouping log messages into sequences based on the block_id present in each log message. Each session is labeled using ground truth. For other datasets, including BGL, Thunderbird, and Liberty, we utilize a sliding window strategy to group log messages, with a window size of 100 messages and a step size of 100 messages. A log sequence is deemed anomalous if it contains at least one anomalous log message according to the ground truth.
@@ V-E Performance Evaluation (RQ1)
Effect of labeled anomalies (RQ2): As illustrated in Table II, in contrast to methods such as DeepLog, LogAnomaly, FastLogAD, LogBERT, and RAPID, which require clean datasets devoid of anomalies to build anomaly detection models, methods like PLELog, LogRobust, CNN, NeuralLog, and LogLLM demonstrate superior performance. These models are trained using not only normal samples but also labeled anomalies. For instance, these five methods achieve an average F1-score above 0.771 across four datasets, whereas others that do not utilize labeled anomalies perform poorly, with an average F1-score below 0.602 across four datasets. This demonstrates that incorporating labeled anomalies can provide a significant advantage to anomaly detection methods.
@@
V-F Effects of Different Preprocessing Techniques (RQ3)
@@
We evaluate the effectiveness of the different preprocessing techniques. The results are shown in Table IV. In this table, ‘Raw’ indicates that the content of log messages is not preprocessed and is directly input into the proposed deep model. ‘Template’ indicates that sequences of log templates produced by Drain [51], a log parser, are used as input for the proposed deep model. ‘Template ID’ signifies that the IDs of log templates, obtained by Drain, are simply encoded into numeric vectors using an embedding layer instead of BERT. The preprocessing technique ‘Template ID’ renders the model unable to capture the semantic information within log messages. Notably, the parser Drain is applied to the entire dataset, rather than only the training dataset, to avoid performance degradation due to the OOV problem. ‘RE’ indicates that regular expressions, as introduced in Section IV-A, are used for preprocessing log messages.
As anticipated, the preprocessing technique ‘RE’ yields the highest F1-score across all datasets. Conversely, the preprocessing technique ‘Template ID’ consistently results in the lowest F1-score across all datasets, averaging 36.8% lower than that of ‘RE’. This can be attributed to the fact that ‘Template ID’ hinders the model’s ability to capture the semantic information within log messages, thereby impairing its capability to detect anomalies from a natural language perspective. The preprocessing techniques ‘Raw’ and ‘Template’ result in relatively good performance, but their F1-scores are still 6.4% and 5.7% lower than that of ‘RE’, respectively. For the preprocessing technique ‘Raw’, the variable parts (parameters that carry dynamic runtime information) within the content of each log message have little influence on anomaly detection. However, due to their high randomness, they can confuse the model, making it difficult to discern anomalies. For the preprocessing technique ‘Template’, the parser is not always reliable, sometimes incorrectly removing the constant parts or retaining the variable parts, which can lead to information loss or confusion for the model, making it difficult to discern anomalies.
V-G Ablation Study of the Training Procedure (RQ4)
@@
We investigate the effect of each training procedure through an ablation study. The results are presented in Table V, where ‘W/O’ denotes ‘without’. We have the following observations:
Skipping any training stage results in a decrease in the F1-score across all datasets, demonstrating the effectiveness of our three-stage training procedure. It is noteworthy that training without stage 1 leads to the worst performance, with the F1-score averaged across all datasets decreasing by as much as 29.7%. However, training without stages 1&2 (only adopting training stage 3: fine-tuning the entire model) yields acceptable performance, with only a 3.2% decrease in the average F1-score. This demonstrates that fine-tuning Llama to capture the answer template (Stage 1) is essential before training the embedder (BERT and projector) of log messages (Stage 2). Without stage 1 (i.e., directly training the embedder), the embedder may be misdirected, resulting in incorrect semantic capture of log messages and model failure. Training without stage 3 yields relatively poor performance, with an average F1-score decrease of 10.5%. This indicates that sequentially fine-tuning Llama and training the embedder alone is insufficient for the model to capture anomalous patterns; cohesive fine-tuning of the entire model is essential. Training without stages 2 and 1&2 also results in a performance decrease, with average F1-score reductions of 2.5% and 3.2%, respectively. This demonstrates that individually training the embedder before fine-tuning the entire model can also enhance performance. This stage allows the embedder to generate better semantic vectors of log messages for Llama to discern anomalies.
@@
V-H Impact of Minority Class Oversampling
Note that normal and anomalous samples in the training dataset are imbalanced, as shown in Table I. For the HDFS, BGL, and Thunderbird datasets, normal samples outnumber anomalous samples. Conversely, in the Liberty dataset, anomalous samples exceed normal samples. As described in Section IV-C1, the hyper-parameter $\beta$ controls the proportion of the minority class by oversampling to address the data imbalance problem. In this section, we investigate the impact of $\beta$ by varying its value. Fig. 4 illustrates the performance of LogLLM on the four datasets under different magnitudes of $\beta$. When $\beta=0$, the samples are not oversampled; instead, the original datasets are utilized directly for training.
@@ V-H Impact of Minority Class Oversampling
As illustrated in Fig. 4c, the trend of the F1-score is basically the same across all datasets. The F1-score increases and then decreases as $\beta$ increases. However, the LogLLM seems not to be sensitive to $\beta$; when $\beta$ is between 10% and 80%, the variation in the F1-score is no more than 0.07. Thanks to the substantial semantic knowledge embedded in LLMs, a trained model can effectively learn anomalous patterns and detect anomalies, even when the minority class constitutes only 10% of the dataset. However, LogLLM appears unable to effectively handle extremely imbalanced scenarios. For instance, in the Thunderbird dataset, anomalies constitute only 1.05% of the samples, causing the trained model to be biased and classify all samples as normal. As a result, precision, recall, and F1-score are all equal to 0.
Compared to the BGL and Thunderbird datasets, the precision, recall and F1-score for the HDFS and Liberty datasets exhibit minimal variation with respect to $\beta$. This consistency arises from the more distinct patterns between abnormal and normal samples in the HDFS and Liberty datasets, allowing LogLLM to easily differentiate them, regardless of the ratio of normal and abnormal samples.